If Ben invests $4000 at 3% interest per year, how much additional money must he invest at 4.5% annual interest to ensure that the interest he receives each year is 3.5% of the total amount invested?
Answer by mananth(16946)

$4000 @ 3 % Interest = $120
..
Let him invest $ x @ 4.5%
interest he will get on this amount = x*0.045*1
=0.045x
..
Total interest will be 120 + 0.045x.............1
....
Total amount invested = 4000+x
interested desired = 3.5%
= 0.035*(4000+x ) ........................2
...
equate 1 &2
120+0.045x =0.035(4000+x)
120+0.045x=140+0.035x
0.045x-0.035x=140-120
0.01x= 20
/0.01
x=20/0.01
x= $2000 . investment @ 4.5%